OpenStudy (anonymous):

PLEASE HELP - What is the distance between points (12, 4) and (18, –1)? Round to the nearest tenth.

OpenStudy (anonymous):

If you think of the distance horizontally and the distance vertically as the sides of a triangle, the third side is the distance you are looking for. So by pythagoras thereom, we know a^2 + b^2 = c^2. a = x2-x1 b= y2 - y1 so c=\[\sqrt{(x2-x1)^{2}+(y2-y1)^{2}}\] and c would be your distance